### v1

# Efficient Optimization Accelerator Framework for Multi-state Ising Problems

Abstract
Ising Machines are a prominent class of hardware architectures that aim to solve NP-Hard combinatorial optimization problems. These machines consist of a network of interacting binary spins/neurons that evolve to represent the optimum ground state energy solution. Generally, combinatorial problems are transformed into quadratic unconstrained binary optimization (QUBO) form to harness the computational efficiency of these Ising machines. However, this transformation, especially for multi-state problems, often leads to a more complex exploration landscape than the original problem, thus severely impacting the solution quality. To address this challenge, we model the spin interactions as a generalized boolean logic function to significantly reduce the exploration space. We benchmark the graph coloring problem from the class of multi-state NP-hard optimization using probabilistic Ising solvers to illustrate the effectiveness of our framework. The proposed methodology achieves similar accuracy compared to state-of-the-art heuristics and machine learning algorithms, and demonstrates significant improvement over the existing Ising methods. Additionally, we demonstrate that combining parallel tempering with our existing framework further reduces the coloring error by up to 50 compared to the conventionally used Gibbs sampling algorithm. We also design a 1024-neuron all-to-all connected probabilistic Ising accelerator that shows up to 10000 performance acceleration compared to heuristics while reducing the number of required physical neurons by 1.5-4 compared to conventional Ising machines. Indeed, this accelerator solution demonstrates improvement across all metrics over the current methods, i.e., energy, performance, area, and solution quality. Thus, this work expands the potential of existing Ising hardware to solve a broad class of these multistate optimization problems.
Department of Electrical Engineering and Computer Sciences, University of California, Berkeley, California 94720, USA
New computing paradigms are getting significant attention due to exponentially growing computing needs [1]. A wide variety of problems falls into the class of non-deterministic polynomial-time hard (NP-hard) and are difficult to solve optimally using conventional computing solutions [2, 3]. The solution space grows exponentially with problem size, therefore brute force searching does not work. However, heuristics [4] and annealing [2] based approaches have been conventionally used to tackle these computationally hard problems in domains such as logistics [5, 6], biology [7], integrated circuits design [8], etc. In this context, Ising machines-based accelerators [9, 10, 11, 12, 13, 14, 15, 16] are currently being leveraged to efficiently find the solution to hard optimization problems. Various technologies and hardware architectures have been explored to build these Ising accelerators including quantum annealing with superconducting qubits [17], classical annealing in memristor or RRAM [18, 19], coherent Ising machines employing optical oscillators [13, 20], coupled oscillators [14, 16, 21], neuromorphic hardware [22] and stochastic circuits (probabilistic bits) [12, 9, 23, 10]. Quantum annealers show promising results [24] but require low temperatures making them expensive in cost and power. Therefore, classical alternatives have gained attention due to their room-temperature operation and realization using the current semiconductor process flow. This work particularly focuses on stochastic/probabilistic Ising machines to efficiently solve combinatorial optimization problems.
Probabilistic Ising architectures follow Boltzmann machine binary neural network principles [25, 26]. These architectures are physically constructed using binary stochastic neurons interacting with each other and aim to minimize the Ising Hamiltonian (Fig. 1a, Supplementary Note 1). The neuron update follows sigmoidal activation (Fig. 1b) such that neurons stochastically move towards the minimum energy states (Fig. 1c). To harness the efficient solution exploration of these hardware architectures, the cost/energy functions of many complex optimization problems are converted into Ising Hamiltonian form. Based on this conversion, these optimization problems can be broadly divided into three categories (Fig. 1d) [27]. The problem class comprised of binary solution state space with no imposed constraints has been widely mapped and efficiently solved on Ising machines [26, 20, 16]. On the other hand, problems with integer/multi-state valued solutions do not naturally convert to binary solution state space, thus leading to additional constraints in the converted Ising Hamiltonian [27, 28, 29]. This direct conversion method frequently causes the Ising machines to explore infeasible solution space, resulting in inefficiencies [18, 30, 31]. One such example is solving the graph coloring problem with Ising machines.
Graph coloring is an NP-hard optimization problem that seeks to assign different colors to the connected nodes of a graph network. It is an example of an integer optimization problem where each integral value represents the color of a node. Previously, annealing and heuristic methods (Tabu-search [32]) were utilized to tackle this problem. These approaches achieve good solution accuracy but suffer from long runtimes for large and densely connected problem instances. Recently, learning-based approaches[33, 34] especially graph neural networks are applied to solve the problem accurately and efficiently at a scale. Conventional Ising hardware or its modified versions are also proposed to solve this problem. However, they significantly lag behind heuristics and learning-based approaches in solution quality. To an extent, earlier works [18, 30, 31] overcome this bottleneck by adopting post-processing techniques using additional hardware and software, which adversely affects area and computation time.
In this work, we present an end-to-end probabilistic Ising implementation that combines advances in multi-state problem mapping, spin interaction design, and an efficient hardware architecture to demonstrate significant improvement in area, solution accuracy, and time-to-solution. First, we propose vectorized mapping that represents the node colors as binary vectors rather than using the customary one-hot form. This circumvents the additional mapping constraint in the Ising Hamiltonian that arises due to one-hot encoding [27]. Thus, it completely discards the exploration of infeasible/invalid solution space and improves the solution quality. Second, the interactions among the binary vector states are modeled using truth tables employed in digital logic. Third, we implement an efficient FPGA accelerator that uses higher-order multiplexers to represent these truth tables. Altogether, this multi-state Ising implementation shows approximately 10000 speed improvement and 5 power improvement compared to GPUs for solving problems up to 256 nodes and 16 colors graph coloring problems. This current methodology is then integrated with the parallel tempering approach to improve the solution quality and provide competitive or even better coloring results than Tabucol heuristics and machine learning algorithms.
Mapping Framework
A wide range of combinatorial optimization problems belongs to multistate/integer-valued solution space, including graph coloring, knapsack problems [35], etc. Graph coloring problem, investigated in this work, has applications, such as layout decomposition[36], register allocation[37], logic minimization[38], scheduling[33], and many more. The problem instance that aims to color nodes with colors, requires physical neurons/nodes in the conventional Ising framework. Because this framework relies on one-hot encoding to represent node colors, additional constraints are imposed in Hamiltonian to enforce the legal combination of one-hot bits [27]. It is written as:
where represents one-hot bit for node . This soft constraint Hamiltonian is expected to be zero for the ground state energy solution. However, it often takes the non-zero value, thereby making Ising samplers inefficient in solving such problems. (see Fig. 2a). The one-hot encoding requires bits to represent possible colors of a node. As a result, Ising machines search over states with only valid states in the solution space. It makes the optimization hard and deteriorates the solution quality. Hence, the conventional Ising framework often fails to solve this class of problems. Despite these challenges, this approach is often adopted for most multistate optimization problems because it leads to quadratic interactions supported in most Ising machines [26, 10, 16, 20]. In probabilistic computing [12], binary multiplexers offer simplistic and efficient realization for state-weight multiplication.
In this work, we propose a vectorized mapping approach (Fig. 2b) to tailor the graph coloring problem using a binary encoding technique. It represents the color state using bits vector and requires physical neurons/nodes for a graph coloring problem with nodes and colors. In this way, it eliminates all invalid state space and removes the extra constraints in Ising Hamiltonian (Eq. 1). Therefore, the proposed approach leads to a solvable energy landscape without any additional complexities. We also adopt parallel tempering [39] to enhance the solution exploration and improve the solution quality.
The proposed vector mapping is a generic approach that can be applied to any multistate optimization problem being solved on Ising accelerators (see Fig. 2b). First, we model the function (see Fig. 2b) in the Hamiltonian for a specific problem. Here multiplication factor enforces edge connectivity. is modeled in a truth table-based format. These truth tables can then be represented via multiplexers to develop a probabilistic accelerator employing the proposed mapping. For graph coloring problems, for the cases when the nodes ( and ) take the same value or color value greater than q (invalid colors), becomes one. is zero for the rest of the cases. The resulting Hamiltonian is a higher-order polynomial. It is minimized when is zero, implying graph coloring criteria are satisfied. The hardware architecture for this scheme is derived from the standard node update rule for probabilistic Ising machines. For bit of node , the update rule is as follows:
where takes binary value 0 and 1. The binary nature of converts the differentiation transforms into equal to - . This is implemented in hardware using higher-order multiplexers followed by a subtraction unit (see Fig. 2), called as vecmul unit. The rest of the hardware implementation remains the same as shown in Fig. 1.
Results
Next we investigate the proposed vectorized mapping approach for the graph coloring problem. For benchmarking, we have used it on publicly available COLOR dataset [40]. First, we implement both probabilistic Ising and proposed vectorized approaches on GPU for comparison. Fig. 3 shows the results for the problem, which is deemed a hard problem [33]. For both schemes, the exploration for solution happens in a similar way (Fig. 3a). For this particular instance, none of the states explored by the Ising framework is valid and therefore unable to find reasonable solutions for most of the problem instances (Fig. 3b) (for more details, see Supplementary section S1). This comparison for other problem instances in the dataset is shown in the Extended Data Fig. S1 and S2. Further, we employ the parallel tempering method to enhance the exploration while sampling the energy landscape. This method involves the simultaneous simulation of M replica systems with vectorized mapping running at different temperature schedules. At high temperatures, the system tends to explore the broad region of the solution space while low temperatures allow precise sampling around a particular region (Fig. 2b). High-temperature replicas ensure that the system doesn’t get stuck around the local minimum solution. The states of higher-temperature replica systems are exchanged with low-temperature systems via a swap rule (see Supplementary Section 2). This state exchange phenomenon enables the exploration of certain parts of solution space which could not be possible via a single temperature schedule. Fig. 3c depicts the energy exploration for the problem using the parallel tempering approach. It clearly shows that the system continuously escapes from the minimum local energy solution and therefore, achieves the best possible solution accuracy.
We benchmark the vectorized mapping solution accuracy for graph coloring problems in the dataset against learning-based approaches including graph neural network (GNNs) [34] and its derivative architectures GraphSage (PI-SAGE) [33] and heuristics-based approach, i.e., Tabucol [32]. GNN-based solvers frame the coloring problem as a multi-class node classification problem and employ unsupervised training by formulating a loss function. In contrast, the Tabucol technique searches the ground state energy solution by moving small steps and maintains a tabu list to avoid cycling around local minima. We implement Tabucol heuristics on Intel Xeon Gold 6330 processor and solve the graph coloring problem instances 200 times with 1000 iteration steps each. For accuracy comparison, we also run Ising and vectorized mapping framework with single-flip Gibbs sampling on Nvidia A100 Tensor Core GPU and solve the same problems 200 times with 1000 iteration steps each. The results comparing the described methods are shown in Table 1. It includes easy, medium, and hard problem instances labeled in work [41]. The Ising approach performs worse by only being able to solve small and easy problem instances. Among GNN-based solvers, the GraphSage architecture (PI-SAGE) offers better solution accuracy but it suffers from longer training times [33]. By contrast, proposed vectorized mapping gives competitive coloring results compared to Tabucol heuristics and PI-SAGE GNN while having slightly lower accuracy for hard-to-solve problem instances. Employing the described parallel tempering with vectorized mapping reduces the error up to 50% on these hard problems and therefore performs better than other methods.
We implement the proposed vectorized architecture (Fig. 2) onto VCU118 FPGA to establish the performance acceleration and energy efficiency benchmarks. The FPGA accelerator uses the memory-mapped IO interface used by software applications to program the problem weights and take the solution out (see Supplementary Section 3 for more details). The accelerator (Fig. 1) fetches the weight data from the memory and does neuron states-weight multiplication. The accumulated product is passed through a sigmoid activation implemented using LUT and further compared with a Linear Feedback Shift Register (LFSR) based random number generator (RNG) to get the updated node value. The architecture runs at 90 MHz clock frequency where only one node gets updated in each cycle. Further, it supports graph coloring problem sizes up to 256 nodes all-to-all connected with a maximum chromatic number of 16, equivalent to 1024 vectorized nodes in total. The framework achieves the same accuracy for dataset problems as the GPU-based vectorized mapping (see Table 1).
The proposed architecture not only gives better accuracy than the baseline Ising-based architecture but also reduces the area implementation. The area implementation is quantified in terms of number of physical nodes required to map any problem instance and solve it. The increase in physical nodes leads to an increase in the number of interactions and hence accumulator size grows leading to a penalty on the implementation area. Fig. 4a shows that the vectorized mapping requires 1.5-4 times fewer nodes for the graph coloring problem instances with chromatic numbers up to 16.
The FPGA implementation of the proposed approach shows 1000 - 10000 speedup compared to Tabucol heuristics implemented on Intel Xeon Gold 6330 CPU cores and its equivalent implementation on Nvidia A100 Tensor Core GPU (see Fig. 4b). Vectorized mapping on GPU could not take advantage of efficient multiplication of binary states with weights and, therefore, suffers with time-to-solution. The same hypothesis will hold for the accelerators that are not built to support these binary calculations. The FPGA accelerator also offers 5 power improvements over GPU implementation. Therefore, accelerating the time-to-solution by 3-4 orders of magnitude on FPGA at a low power budget significantly improves energy efficiency.
Conclusion
Current methods employ the brute-force framework to map and solve multistate Ising problems onto conventional Ising machines by using one-hot state encoding. This work presents an alternative approach of vectorized mapping that maps the state using binary encoding. It not only reduces state exploration from to for a q color problem instance but also removes the constraints resulting from the conventional brute-force Ising approach. As a result, the proposed approach converges close to the optimal solution achieved by heuristics and learning based approaches. Exising work [16, 12, 20] on Ising Machines and its derivative frameworks generally aims to achieve accuracy within a certain error range compared to heuristics. However, we show that our method combined with parallel tempering achieves the solution accuracy even better than the heuristics for some of the problems, while retaining the ability to significant speed up on a custom accelerator. We also present a generalized truth table-based method that can be leveraged to map other multi-state problems. Using probabilistic Ising machine architectures, these truth tables to model the neuron interactions can be directly mapped onto higher order multiplexers. We have implemented this architecture on FPGA to benchmark time-to-solution, energy efficiency, and area efficiency. Overall, the hardware implementation consumes 5W power and achieves approximately 1000 - 10000 speedup compared to heuristics on CPU and vectorized mapping on GPU. The presented hardware and software framework provides a new way to substantially expand the capabilities of the Ising machines to accurately handle a wide range of multistate optimization problems in a performance and energy-efficient manner.
1. Probabilistic Ising Machines
Probabilistic Ising machines follow the principle of the Boltzmann Machine binary neural network. Therefore, the probability distribution corresponding to a given state where is the normalizing/partition function. The distribution is sampled such that the Ising machine stochastically moves towards higher probability states and lower energy. It is enabled by appropriately formulating the update rule which for Hamiltonian of the form , is given by where .
2. Parallel Tempering
The parallel tempering algorithm utilizes multiple Markov chains (replica chains) running at different temperature schedules representing different probability distributions. It allows a broader exploration of the energy landscape and facilitates better mixing to avoid local energy minima solutions. In this work, we implement 100 parallel chains running at a constant temperature geometrically spaced from temp0 (0.01) and tempn (40). After every 15 sampling steps, adjacent pairs of chains are swapped alternatively staring with odd and even chain indexes. The update rule is given as:
3. FPGA Setup PCIe
In this work, we implement the vectorized mapping on the Xilinx Virte UltraScale+ VCU118 FPGA evaluation kit. The FPGA is interfaced with CPU using peripheral component interconnect express (PCIe) interface. In particular, we use open-source xillybus IP core for the interface with data transfer capabilities of 50MB/s. The data is transferred via a memory mapped interface implemented using block memory and a designed memory controller.
The digital implementation of 256 nodes and 16 color accelerators supports a network of 1024 probabilistic Ising nodes. These nodes consist of an LFSR-based pseudorandom number generator with a fixed seed, a lookup table-based sigmoidal activation, higher order multiplexer-based matrix multiplication, and a threshold that generates the output state of the neuron. These states are stored in the local memory of FPGA and then transfered to CPU via PCIe interface.
This work is supported by MURI project from Department of Defense and Office of Naval Research.
C.G. formulated the vectorized scheme, performed the CPU/GPU benchmarks and designed the vectorized accelerator on FPGA. C.G, and S.S co-wrote the manuscript; S.S supervised the research. All authors contributed to discussions and commented on the manuscript.
The authors declare that they have no competing financial interests.
Correspondence and requests for materials can be addressed to either C.G.
(chiraggarg@berkeley.edu) or S.S. (sayeef@berkeley.edu).
Code and Data will be made available on reasonable request by emailing C.G. or S.S.
a, Quadratic unconstrained binary optimization (QUBO) problems are represented as an Ising network with nodes representing the states and edges depicting interactions between them. b, Probabilistic Ising architecture based on the Boltzmann machine update rule to minimize the problem Hamiltonian and energy function. c, The State of the Ising machine evolves and converges to the optimal energy solution. d, Classification of optimization problems based on the state values and problem constraints.
a, The baseline Ising framework employs one-hot encoding to map node color graph coloring problem to nodes. The encoding constraint is enforced by adding an extra term in the Hamiltonian leading to a complex optimization landscape. Since only states are valid among states for a node, this mapping often explores a lot of invalid solution space making it hard for Ising hardware to reach the optimal energy solution. This mapping preserves QUBO Hamiltonian, therefore, the interactions are modeled via binary multiplexers in hardware. b, Vectorized framework maps the color state of a node using hardware nodes. It allows exploration only in valid solution space making it easier for hardware to reach ground-state energy solutions. The solution exploration is further enhanced by the adoption of parallel tempering. This mapping can be modeled in truth table format which is directly mapped to a higher-order multiplexer in hardware.
a, Energy evolution for problem instance mapped and solved using Ising and Vectorized mapping implementation on GPU. (b) Valid and Invalid energy states recorded during solution exploration employing Ising and Vectorized implementation. Ising framework explores completely over invalid solution space making it hard to reach an optimal solution. (c) Vectorized framework combined with parallel tempering enhances the solution exploration avoiding the states being stuck around local minima solution.
a, Physical implementation nodes used in Ising and Vectorized mapping approach for dataset problem [40]. (b) Time-to-solution (TTS) comparison of vectorized mapping implemented on FPGA with Tabucol heuristic on CPU and it equivalent implementation on GPU.
Problem nodes edges colors GNN[33] PI-SAGE[33] Tabucol Probabilistic Ising Vectorized GPU Vectorized FPGA Vectorized+ Parallel Tempering GPU anna 138 493 11 1 0 0 - 0 0 0 myciel5 47 236 6 0 0 0 0 0 0 0 myciel6 95 755 7 0 0 0 0 0 0 0 queen55 25 160 5 0 0 0 0 0 0 0 queen66 36 290 7 4 0 0 - 1 1 0 queen77 49 476 7 15 0 0 - 6 5 0 queen88 64 728 9 7 1 0 - 4 2 1 queen99 81 1056 10 13 1 0 - 5 4 2 queen812 96 1368 12 7 0 0 - 2 2 0 queen1111 121 1980 11 33 17 15 - 20 18 14 queen1313 169 3328 13 40 26 21 - 31 26 21